The demand and supply equations for a certain item are given by
D=-5p+40
S=-p^2+30p-8
Find the equilibrium price.

Answer by stanbon(75887)

D=-5p+40
S=-p^2+30p-8
Find the equilibrium price.
--------------------------------------
-5p+40 = -p^2+30p-8
p^2-35p+48=0
p = [35 +- sqrt(35^2-4*48)]/2
p = [35 +- sqrt(1033)]/2
p = [35 +- 32.14]/2
p = 33.57 or p = 1.43
========================
Cheers,
Stan H.